OpenStudy (babyjayy):

The minimum and maximum temperature on a cold day in Lollypop Town can be modeled by the equation below: 2|x − 6| + 14 = 38 What are the minimum and maximum temperatures for this day? x = −9, x = 21 x = −6, x = 18 x = 6, x = −18 No solution its either no solution or b

OpenStudy (zin7770):

x=-6, x=18 The answer is B. I am 99% sure.

OpenStudy (imstuck):

Zin7770 is correct. It's B.

OpenStudy (imstuck):

I'm 100% sure! ; )

OpenStudy (zzr0ck3r):

I am 100% sure you should all read the rules

OpenStudy (acespeedfighter):

███████████▓▓▓▓▓▓▓▓▒░░░░░▒▒░░░░░░░▓█████ ██████████▓▓▓▓▓▓▓▓▒░░░░░▒▒▒░░░░░░░░▓████ █████████▓▓▓▓▓▓▓▓▒░░░░░░▒▒▒░░░░░░░░░▓███ ████████▓▓▓▓▓▓▓▓▒░░░░░░░▒▒▒░░░░░░░░░░███ ███████▓▓▓▓▓▓▓▓▒░░▒▓░░░░░░░░░░░░░░░░░███ ██████▓▓▓▓▓▓▓▓▒░▓████░░░░░▒▓░░░░░░░░░███ █████▓▒▓▓▓▓▓▒░▒█████▓░░░░▓██▓░░░░░░░▒███ ████▓▒▓▒▒▒░░▒███████░░░░▒████░░░░░░░░███ ███▓▒▒▒░░▒▓████████▒░░░░▓████▒░░░░░░▒███ ██▓▒▒░░▒██████████▓░░░░░▓█████░░░░░░░███ ██▓▒░░███████████▓░░░░░░▒█████▓░░░░░░███ ██▓▒░▒██████████▓▒▒▒░░░░░██████▒░░░░░▓██ ██▓▒░░▒███████▓▒▒▒▒▒░░░░░▓██████▓░░░░▒██ ███▒░░░░▒▒▒▒▒▒▒▒▒▒▒▒░░░░░░███████▓░░░▓██ ███▓░░░░░▒▒▒▓▓▒▒▒▒░░░░░░░░░██████▓░░░███ ████▓▒▒▒▒▓▓▓▓▓▓▒▒▓██▒░░░░░░░▓███▓░░░░███ ██████████▓▓▓▓▒▒█████▓░░░░░░░░░░░░░░████ █████████▓▓▓▓▒▒░▓█▓▓██░░░░░░░░░░░░░█████ ███████▓▓▓▓▓▒▒▒░░░░░░▒░░░░░░░░░░░░██████ ██████▓▓▓▓▓▓▒▒░░░░░░░░░░░░░░░░▒▓████████ ██████▓▓▓▓▓▒▒▒░░░░░░░░░░░░░░░▓██████████ ██████▓▓▓▓▒▒██████▒░░░░░░░░░▓███████████ ██████▓▓▓▒▒█████████▒░░░░░░▓████████████ ██████▓▓▒▒███████████░░░░░▒█████████████ ██████▓▓░████████████░░░░▒██████████████ ██████▓░░████████████░░░░███████████████ ██████▓░▓███████████▒░░░████████████████ ██████▓░███████████▓░░░█████████████████ ██████▓░███████████░░░██████████████████ ██████▓▒██████████░░░███████████████████ ██████▒▒█████████▒░▓████████████████████ ██████░▒████████▓░██████████████████████ ██████░▓████████░███████████████████████ ██████░████████░▒███████████████████████ █████▓░███████▒░████████████████████████ █████▒░███████░▓████████████████████████ █████░▒██████░░█████████████████████████ █████░▒█████▓░██████████████████████████ █████░▓█████░▒██████████████████████████ █████░▓████▒░███████████████████████████ █████░▓███▓░▓███████████████████████████ ██████░▓▓▒░▓████████████████████████████ ███████▒░▒██████████████████████████████ ████████████████████████████████████████ ████████████████████████████████████████

OpenStudy (mathstudent55):

You need to solve the absolute value equation you were given. First, you need to isolate the absolute value part. 2|x − 6| + 14 = 38 Subtract 14 from both sides, then divide both sides by 2. 2|x − 6| + 14 - 14 = 38 - 14 2|x - 6| = 24 \(\dfrac{2|x - 6|}{2} = \dfrac{24}{2} \) |x - 6| = 12 We have the absolute value part isolated on the left side. Now we can deal with solving the absolute value equation. Here is the rule: ************************************************************************* To solve the absolute value equation |X| = k, where X is an expression in x and k is a non-negative number, solve the compound equation below: X = k or X = -k *************************************************************************

OpenStudy (mathstudent55):

In our case we have |x - 6| = 12 For us, the expression in x is x - 6. The non-negative number is 12. That means the absolute value equation turns into this compound equation: x - 6 = 12 or x - 6 = -12 We add 6 to both sides of both equations to get x - 6 + 6 = 12 + 6 or x - 6 + 6 = -12 + 6 x = 18 or x = -6